(5+x)g of Hydrogen, Ymol. Of Neon, Pmol. Of Argon and 20g of Helium were mixed together and the total pressure Exerted by the mixture was found to be 800mmHg. Given that the partial pressure of Hydogen, Helium, Argon and Ne are ZmmHg, 1000mmHg, QmmHg and 2400mmHg respectively, determine the value of Y, P, Z and Q


1
Expert's answer
2021-09-23T02:37:01-0400

1000/2400

0.417

Hydrogen = 5+0.417 = 5.147

Helium = 20×0.417 = 8.33